# Helicity controls the direction of fluxes in rotating turbulence

### v1

@@ Appendix B Mean-wave kinetic theory
$\partial_{t}(spe_{\bm{p}}^{s})+\partial_{p_{y}}(sp\Pi^{s})=-\mathbb{1}_{\bm{p}\in A}sp~\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\bm{p}\bm{k}}^{s,-s}+sp\epsilon\chi_{\bm{p}}^{s}.$ (50)
Due to scale separation $k_{y}/p\ll 1$, the helicity transfer between the waves and the mean flow can be neglected. Indeed, decomposing the condensate mode $\bm{k}$ into its two chiralities, the helicity transfer with the waves is
$\mathcal{H}_{\bm{k}}=k(\mathcal{E}_{\bm{k}}^{+}-\mathcal{E}_{\bm{k}}^{-})$ (51)
with $\mathcal{E}_{\bm{k}}^{\pm}$ the energy transfer between the waves and the 2D mode of polarity $\pm 1$. $\mathcal{E}_{\bm{k}}^{\pm}$ can be explicitly written from the helical decomposition of the wave field, but this does not matter for the argument. Scale selection $k_{y}/k_{f}\ll 1$ implies that the 2D-3D helicity exchange (51) is negligible: at most, $\mathcal{H}_{\bm{k}}=O(k_{y}\epsilon)$, in the (unexpected) case where one mean-flow polarity absorbs all the wave energy, hence $\mathcal{H}_{\bm{k}}\ll k_{f}\epsilon^{s}$, the amount of helicity injected in each wave sector $s$. Therefore, the waves conserve their helicity separately, independently of the type of condensate-wave interaction (homochiral or heterochiral). Note that conservation of total helicity of the fluctuations implies that
$\langle uv\rangle_{\bm{p}\bm{k}}^{s,-s}=\langle uv\rangle_{\bm{p}\bm{k}}^{-s,s}$ (52)
and we see that the first term on the RHS of Eq. (50) corresponds to the change in helicity of $\text{sign}(s)$ due to interactions with modes of opposite helicity, mediated by the condensate. In particular, if there is a finite (and positive) energy flux at arbitrary large scales $p$, $\lim_{p\to\infty}\Pi^{s}_{\text{adv}}>0$, then this flux serves as the source for the generation of helicity of sign $s$. The helicity flux $sp\Pi^{s}$ must therefore increase with $p$ and be of sign $s$. From Eq. (50), this implies that $\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\bm{p}\bm{k}}^{s,-s}<0$ for large enough $p$, which in turn implies that the cross-helicity Reynolds stress acts on the mean flow to extract energy from it.
@@ Appendix C Homochiral-wave interactions
Here we focus on Sector H, where heterochiral interactions are off-resonant, hence $\langle a_{\bm{p}}^{s}a_{-\bm{p}+\bm{k}}^{-s}\rangle=0$. The procedure is exactly that in [gome2025waves]. Due to the presence of two sign-definite invariants in this sector (energy and single-sign helicity $H^{s}\equiv\sum spe_{\bm{p}}^{s}$, the strady-state solution is straightforward. Equation (40) for $\bm{p}\in H$ can be written in steady state in the form
$\partial_{p_{y}}\left(-sp\frac{U^{\prime}p_{x}}{\sqrt{2}}\Phi^{s}_{\bm{p}\bm{k}}\right)\equiv\partial_{p_{y}}\left(\Pi_{H^{s}}\right)=sp\epsilon\chi_{p}^{s}$ (53)
@@
$\Phi_{\bm{p}\bm{k}}^{s}$ $=-\frac{\sqrt{2}\epsilon}{{U^{\prime}}p_{x}}\int_{-\infty}^{p_{y}}\frac{q}{p}~\chi^{s}_{\bm{q}}~dq_{y}~~\text{ for }{U^{\prime}}p_{x}<0~(dq_{y}>0);$
$\Phi_{\bm{p}\bm{k}}^{s}$ $=-\frac{\sqrt{2}\epsilon}{{U^{\prime}}p_{x}}\int_{\infty}^{p_{y}}\frac{q}{p}~\chi^{s}_{\bm{q}}~dq_{y}~~\text{ for }{U^{\prime}}p_{x}>0~(dq_{y}<0).$ (54)
@@
Solution (54) is simplified when $\chi_{\bm{p}}^{s}=\chi_{\bm{p}}^{-s}=\delta(p-k_{f})/(8\pi k_{f}^{2}\mathcal{V})$ and leads to the small-scale energy flux:
$\Pi^{s}_{\bm{p}}\equiv\frac{-{U^{\prime}}p_{x}}{\sqrt{2}}\Phi_{\bm{p}\bm{k}}^{s}=\left\{\begin{aligned} &~0~~\text{ if }\sigma p_{y}<-p_{y}^{f}\\ &\frac{\sigma\epsilon^{s}_{p_{x},p_{z}}}{2}\frac{k_{f}}{p}~~\text{ if }-p_{y}^{f}<\sigma p_{y}<p_{y}^{f}\\ &\sigma\epsilon^{s}_{p_{x},p_{z}}\frac{k_{f}}{p}~~\text{ if }\sigma p_{y}>p_{y}^{f}\end{aligned}\right$. (57)
@@
$\begin{split}\epsilon^{s}_{p_{x},p_{z}}&\equiv\int_{-\infty}^{\infty}\chi^{s}_{\bm{q}}~F_{\bm{q}-\bm{k}-\bm{q}}~dq_{y}=\frac{1}{8\pi k_{f}^{2}\mathcal{V}}\int_{-\infty}^{\infty}\delta\left(q-k_{f}\right)F_{\bm{p}\bm{q}}~dq_{y}\\ &=\frac{1}{8\pi k_{f}^{2}\mathcal{V}}\frac{2k_{f}}{p_{y}}F_{k_{f}}=\frac{\epsilon F_{k_{f}}}{4\pi k_{f}\mathcal{V}\sqrt{k_{f}^{2}-p_{x}^{2}-p_{z}^{2}}}\end{split}$ (58)
@@
$F_{k_{f}}\equiv F_{\bm{p},-\bm{k}-\bm{k}}\delta_{p-k_{f}}$ (59)
equal to the resonant filter (34) evaluated at the forcing scale. Here we have multiplied the energy injection rate by the filter to take into account the fact that some of the modes decouple from the condensate, if $\Delta^{ss}\approx 0\Leftrightarrow\omega_{\bm{p}\bm{q}}^{ss}\delta_{p-k_{f}}<{U^{\prime}}$ at the forcing scale (which implies that $\Delta_{\bm{p}}^{s,-s}\approx 0$). Such modes only interact with other 3D waves are then treated as if they are unforced in the QL system. Then, $\Phi_{\bm{p}\bm{k}}^{s}=0$ and the corresponding Reynolds stress is zero. The decoupling of 2D-3D interactions as $\Omega\to\infty$ is discussed in [gome2025waves].
Solution (57) reflects the conservation of helicity within each polar sector: $p\Pi^{s}=k_{f}\epsilon_{p_{x},p_{z}}^{s}$ with the energy flux $\Pi^{s}=-U^{\prime}p_{x}\Phi_{\bm{p}\bm{k}}^{s}$. This shows that the conservation of helicity by sign imposes a decaying energy flux to small scales, $\Pi^{s}\sim k_{f}/p\epsilon_{p_{x},p_{z}}^{s}$. The sign of $\Pi^{s}$ depends on the sign of $U^{\prime}p_{x}p_{y}$, as follows from the choice of boundary conditions (56): it is positive when $p_{y}>0$ and ${U^{\prime}}p_{x}<0$.
@@
$\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\bm{p}\bm{k}}$ $=-\frac{{U^{\prime}}}{\sqrt{2}}\frac{p_{x}p_{y}}{p^{2}}(\Phi_{\bm{p}\bm{k}}^{+}+\Phi_{\bm{p}\bm{k}}^{-})$ (60)
@@
the total energy injected in the $(p_{x},p_{z})$ layer. This solution is shown in Fig. 5. Importantly, $\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\bm{p}\bm{k}}$ is positive when $\sigma p_{y}>0$: energy goes from the waves to the 2D condensate.
Figure 5: Reynolds stress in the spectral plane ($p_{y},p_{z},p_{x}=8\pi/L_{x}$), computed when interactions are homochiral (Sector H). This corresponds to the high-rotation regime. Advection by the shear flow generates a decaying forward flux of wave energy, which is transferred to the shear flow (${U^{\prime}}\langle uv\rangle_{\bm{p}\bm{k}}>0$).
@@ Appendix D Heterochiral interactions: governing equations
System (40)-(41) when all modes $\bm{p}$ are in Sector A describes rotation-free interactions between 3D modes, excited at scale $l_{f}$, and an imposed large-scale mean flow $\bm{U}(y)$. This is due to the frequencies of 3D modes in this sector being too slow, so that $\Delta_{\bm{p}}^{+-}\approx 1$ and rotation has no effect on the QL dynamics at leading order. For these modes, the same dynamical system would be obtained if $\Omega$ were put to 0 in Eq. (27).
@@
However, the helical basis does not diagonalize the stationary system of correlators $\Phi_{\bm{p}}^{s},\Psi_{\bm{p}}^{+-}$. Instead, the system can be triangularized, and possibly solved analytically, when formulating the fluid equations in terms of primary variables, as is classically done in mean-fluctuation descriptions of shear flows [schmid2012stability, farrell1993optimal]. In this section we will present the solution for the energy flux and Reynolds stress for this non-rotating, directly forced case. This solution will be useful for the treatment of modes in Sector A, which also exhibit such dynamics and can either be directly forced, or receive their energy through a flux in $p$ space, coming from modes in the homochiral sector.
@@
We consider a non-rotating system of mean flow $\bm{U}=U(y)\mathbf{e}_{x}$ with a given mean shear rate $U^{\prime}$}, and 3D perturbations $[u,v,w](x,y,z)$ such that $\bm{u}\cdot\nabla\bm{u}\ll\bm{u}\cdot\nabla\bm{U}$, which obeys the Rayleigh equations
$\partial_{t}\begin{bmatrix}\nabla^{2}v\\ \eta\end{bmatrix}=\begin{bmatrix}-(U\nabla^{2}-\partial_{y}^{2}U)\partial_{x}&0\\ -{U^{\prime}}\partial_{z}&-U\partial_{x}\end{bmatrix}\begin{bmatrix}v\\ \eta\end{bmatrix}+\begin{bmatrix}\nabla^{2}f_{y}\\ (\nabla\times\bm{f})_{y}\end{bmatrix}$ (63)
@@
with $\eta\equiv(\nabla\times\bm{u})_{y}$ the $y$-component of vorticity (see e.g [farrell1993optimal]). Restricting Eq. (63) to our sinusoidal mean flow $U(y)={U^{\prime}}L_{y}/(\sqrt{2}\pi)\sin(2\pi y/L_{y}),U_{\bm{k}}=\frac{U^{\prime}}{i\sqrt{2}k_{y}}$ and using scale separation $L_{y}k_{f}\gg 1$ (consequently, $\partial_{y}^{2}U\ll U\nabla^{2}$), we obtain the simplified system in Fourier space:
$\partial_{t}(p^{2}v_{\bm{p}})$ $=\frac{p_{x}{U^{\prime}}}{\sqrt{2}k_{y}}\Bigg[|-\bm{p}-\bm{k}|^{2}v_{-\bm{p}-\bm{k}}^{*}-|-\bm{p}+\bm{k}|^{2}v_{-\bm{p}+\bm{k}}^{*}\Bigg]+p^{2}f_{\bm{p},y}$ (64)
$\partial_{t}\eta_{\bm{p}}$ $=\frac{{U^{\prime}}}{\sqrt{2}}\Bigg[\frac{p_{x}}{k_{y}}(\eta_{-\bm{p}-\bm{k}}^{*}-\eta_{-\bm{p}+\bm{k}}^{*})-ip_{z}(v_{-\bm{p}-\bm{k}}^{*}+v_{-\bm{p}+\bm{k}}^{*})\Bigg]+g_{y}$ (65)
with $k_{y}=2\pi/L_{y}$ and $g_{y}=ip_{z}f_{\bm{p},x}-ip_{x}f_{\bm{p},z}$ and where we have used that $p_{x}=-q_{x},p_{y}=-q_{y}\pm k_{y},p_{z}=-q_{z}$.
@@
We now write the equation for the ensemble-averaged cross-flow energy $\langle|v_{\bm{p}}|^{2}\rangle$, velocity-vorticity correlation $\langle\eta_{\bm{p}}v_{-\bm{p}}\rangle$ and cross-flow enstrophy $\langle|\eta_{\bm{p}}|^{2}\rangle$ at $O(k_{y}/p)$:
@@
$\partial_{t}\frac{\langle|v_{\bm{p}}|^{2}\rangle}{2}$ $=\frac{{U^{\prime}}}{\sqrt{2}}\Big[p_{x}\partial_{p_{y}}+\frac{4p_{x}p_{y}}{p^{2}}\Big]\langle vv\rangle_{\bm{p}\bm{k}}+\epsilon\chi_{\bm{p},y}$ (66)
@@
$\partial_{t}\frac{\langle\eta_{\bm{p}}v_{-\bm{p}}\rangle}{2}$ $=\frac{{U^{\prime}}}{\sqrt{2}}\Bigg(\Big[p_{x}\partial_{p_{y}}+2\frac{p_{x}p_{y}}{p^{2}}\Big]\langle\eta v\rangle_{\bm{p}\bm{k}}-ip_{z}\langle vv\rangle_{\bm{p}\bm{k}}\Bigg)$ (67)
@@
$\partial_{t}\frac{\langle|\eta_{\bm{p}}|^{2}\rangle}{2}$ $=\frac{{U^{\prime}}}{\sqrt{2}}\Big(p_{x}\partial_{p_{y}}\langle\eta\eta\rangle_{\bm{p}\bm{k}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-}2p_{z}\Im\langle\eta v\rangle_{\bm{p}\bm{k}}\Big)+\epsilon p^{2}\chi_{\bm{p},y}$ (68)
@@
$\langle vv\rangle_{\bm{p}\bm{k}}\equiv\frac{1}{2}\Big(\langle v_{\bm{p}}v_{-\bm{p}+\bm{k}}\rangle+\langle v_{\bm{p}}v_{-\bm{p}-\bm{k}}\rangle\Big),~~~~~\langle\eta v\rangle_{\bm{p}\bm{k}}\equiv\frac{1}{2}\Big(\langle\eta_{\bm{p}}v_{-\bm{p}+\bm{k}}\rangle+\langle\eta_{\bm{p}}v_{-\bm{p}-\bm{k}}\rangle\Big),~~~~~\langle\eta\eta\rangle_{\bm{p}\bm{k}}\equiv\frac{1}{2}\Big(\langle\eta_{\bm{p}}\eta_{-\bm{p}+\bm{k}}\rangle+\langle\eta_{\bm{p}}\eta_{-\bm{p}-\bm{k}}\rangle\Big)$ (69)
and used that $\langle vv\rangle_{\bm{p}\bm{k}}^{*}=\langle vv\rangle_{\bm{p}\bm{k}}$ since $\langle v_{\bm{p}}v_{-\bm{p}+\bm{k}}\rangle=\langle v_{-\bm{p}}v_{\bm{p}+\bm{k}}\rangle+O(k_{y}/p)$, as well as
$\langle v_{\bm{p}}v_{-\bm{p}-\bm{k}}\rangle$ $=\langle v_{-\bm{p}}v_{\bm{p}-\bm{k}}\rangle+O(k_{y}/p),~~~~~~~\langle v_{\bm{p}}v_{-\bm{p}-\bm{k}}\rangle=\langle v_{-\bm{p}}v_{\bm{p}-\bm{k}}\rangle+O(k_{y}/p),~~~~~~~\langle\eta_{\bm{p}}v_{-\bm{p}+\bm{k}}\rangle=\langle\eta_{\bm{p}-\bm{k}}v_{-\bm{p}}\rangle+O(k_{y}/p),$
$k_{y}\partial_{p_{y}}\langle\eta v\rangle_{\bm{p}\bm{k}}$ $=\langle\eta_{p+k}v_{-p}\rangle-\langle\eta_{p}v_{-p+k}\rangle+O(k_{y}/p)=\langle\eta_{p}v_{-p-k}\rangle-\langle\eta_{p-k}v_{-p}\rangle+O(k_{y}/p),$ (70)

### v2

@@ Appendix B Mean-wave kinetic theory
$\partial_{t}(spe_{\boldsymbol{p}}^{s})+\partial_{p_{y}}(sp\Pi^{s})=-\mathbb{1}_{\boldsymbol{p}\in A}sp~\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{s,-s}+sp\epsilon\chi_{\boldsymbol{p}}^{s}.$ (50)
Due to scale separation $k_{y}/p\ll 1$, the helicity transfer between the waves and the mean flow can be neglected. Indeed, decomposing the condensate mode $\boldsymbol{k}$ into its two chiralities, the helicity transfer with the waves is
$\mathcal{H}_{\boldsymbol{k}}=k(\mathcal{E}_{\boldsymbol{k}}^{+}-\mathcal{E}_{\boldsymbol{k}}^{-})$ (51)
with $\mathcal{E}_{\boldsymbol{k}}^{\pm}$ the energy transfer between the waves and the 2D mode of polarity $\pm 1$. $\mathcal{E}_{\boldsymbol{k}}^{\pm}$ can be explicitly written from the helical decomposition of the wave field, but this does not matter for the argument. Scale selection $k_{y}/k_{f}\ll 1$ implies that the 2D-3D helicity exchange (51) is negligible: at most, $\mathcal{H}_{\boldsymbol{k}}=O(k_{y}\epsilon)$, in the (unexpected) case where one mean-flow polarity absorbs all the wave energy, hence $\mathcal{H}_{\boldsymbol{k}}\ll k_{f}\epsilon^{s}$, the amount of helicity injected in each wave sector $s$. Therefore, the waves conserve their helicity separately, independently of the type of condensate-wave interaction (homochiral or heterochiral). Note that conservation of total helicity of the fluctuations implies that
$\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{s,-s}=\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{-s,s}$ (52)
and we see that the first term on the RHS of Eq. (50) corresponds to the change in helicity of $\text{sign}(s)$ due to interactions with modes of opposite helicity, mediated by the condensate. In particular, if there is a finite (and positive) energy flux at arbitrary large scales $p$, $\lim_{p\to\infty}\Pi^{s}_{\text{adv}}>0$, then this flux serves as the source for the generation of helicity of sign $s$. The helicity flux $sp\Pi^{s}$ must therefore increase with $p$ and be of sign $s$. From Eq. (50), this implies that $\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{s,-s}<0$ for large enough $p$, which in turn implies that the cross-helicity Reynolds stress acts on the mean flow to extract energy from it.
@@ Appendix C Homochiral-wave interactions
Here we focus on Sector H, where heterochiral interactions are off-resonant, hence $\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}+\boldsymbol{k}}^{-s}\rangle=0$. The procedure is exactly that in [26]. Due to the presence of two sign-definite invariants in this sector (energy and single-sign helicity $H^{s}\equiv\sum spe_{\boldsymbol{p}}^{s}$, the strady-state solution is straightforward. Equation (40) for $\boldsymbol{p}\in H$ can be written in steady state in the form
$\partial_{p_{y}}\left(-sp\frac{U^{\prime}p_{x}}{\sqrt{2}}\Phi^{s}_{\boldsymbol{p}\boldsymbol{k}}\right)\equiv\partial_{p_{y}}\left(\Pi_{H^{s}}\right)=sp\epsilon\chi_{p}^{s}$ (53)
@@
$\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s}$ $=-\frac{\sqrt{2}\epsilon}{{U^{\prime}}p_{x}}\int_{-\infty}^{p_{y}}\frac{q}{p}~\chi^{s}_{\boldsymbol{q}}~dq_{y}~~\text{ for }{U^{\prime}}p_{x}<0~(dq_{y}>0);$
$\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s}$ $=-\frac{\sqrt{2}\epsilon}{{U^{\prime}}p_{x}}\int_{\infty}^{p_{y}}\frac{q}{p}~\chi^{s}_{\boldsymbol{q}}~dq_{y}~~\text{ for }{U^{\prime}}p_{x}>0~(dq_{y}<0).$ (54)
@@
Solution (54) is simplified when $\chi_{\boldsymbol{p}}^{s}=\chi_{\boldsymbol{p}}^{-s}=\delta(p-k_{f})/(8\pi k_{f}^{2}\mathcal{V})$ and leads to the small-scale energy flux:
$\Pi^{s}_{\boldsymbol{p}}\equiv\frac{-{U^{\prime}}p_{x}}{\sqrt{2}}\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s}=\left\{\begin{aligned} &~0~~\text{ if }\sigma p_{y}<-p_{y}^{f}\\ &\frac{\sigma\epsilon^{s}_{p_{x},p_{z}}}{2}\frac{k_{f}}{p}~~\text{ if }-p_{y}^{f}<\sigma p_{y}<p_{y}^{f}\\ &\sigma\epsilon^{s}_{p_{x},p_{z}}\frac{k_{f}}{p}~~\text{ if }\sigma p_{y}>p_{y}^{f}\end{aligned}\right$. (57)
@@
$\begin{split}\epsilon^{s}_{p_{x},p_{z}}&\equiv\int_{-\infty}^{\infty}\chi^{s}_{\boldsymbol{q}}~F_{\boldsymbol{q}-\boldsymbol{k}-\boldsymbol{q}}~dq_{y}=\frac{1}{8\pi k_{f}^{2}\mathcal{V}}\int_{-\infty}^{\infty}\delta\left(q-k_{f}\right)F_{\boldsymbol{p}\boldsymbol{q}}~dq_{y}\\ &=\frac{1}{8\pi k_{f}^{2}\mathcal{V}}\frac{2k_{f}}{p_{y}}F_{k_{f}}=\frac{\epsilon F_{k_{f}}}{4\pi k_{f}\mathcal{V}\sqrt{k_{f}^{2}-p_{x}^{2}-p_{z}^{2}}}\end{split}$ (58)
@@
$F_{k_{f}}\equiv F_{\boldsymbol{p},-\boldsymbol{k}-\boldsymbol{k}}\delta_{p-k_{f}}$ (59)
equal to the resonant filter (34) evaluated at the forcing scale. Here we have multiplied the energy injection rate by the filter to take into account the fact that some of the modes decouple from the condensate, if $\Delta^{ss}\approx 0\Leftrightarrow\omega_{\boldsymbol{p}\boldsymbol{q}}^{ss}\delta_{p-k_{f}}<{U^{\prime}}$ at the forcing scale (which implies that $\Delta_{\boldsymbol{p}}^{s,-s}\approx 0$). Such modes only interact with other 3D waves are then treated as if they are unforced in the QL system. Then, $\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s}=0$ and the corresponding Reynolds stress is zero. The decoupling of 2D-3D interactions as $\Omega\to\infty$ is discussed in [26].
Solution (57) reflects the conservation of helicity within each polar sector: $p\Pi^{s}=k_{f}\epsilon_{p_{x},p_{z}}^{s}$ with the energy flux $\Pi^{s}=-U^{\prime}p_{x}\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s}$. This shows that the conservation of helicity by sign imposes a decaying energy flux to small scales, $\Pi^{s}\sim k_{f}/p\epsilon_{p_{x},p_{z}}^{s}$. The sign of $\Pi^{s}$ depends on the sign of $U^{\prime}p_{x}p_{y}$, as follows from the choice of boundary conditions (56): it is positive when $p_{y}>0$ and ${U^{\prime}}p_{x}<0$.
@@
$\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ $=-\frac{{U^{\prime}}}{\sqrt{2}}\frac{p_{x}p_{y}}{p^{2}}(\Phi_{\boldsymbol{p}\boldsymbol{k}}^{+}+\Phi_{\boldsymbol{p}\boldsymbol{k}}^{-})$ (60)
@@
the total energy injected in the $(p_{x},p_{z})$ layer. This solution is shown in Fig. 5. Importantly, $\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ is positive when $\sigma p_{y}>0$: energy goes from the waves to the 2D condensate.
Figure 5: Reynolds stress in the spectral plane ($p_{y},p_{z},p_{x}=8\pi/L_{x}$), computed when interactions are homochiral (Sector H). This corresponds to the high-rotation regime. Advection by the shear flow generates a decaying forward flux of wave energy, which is transferred to the shear flow (${U^{\prime}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}>0$).
@@ Appendix D Heterochiral interactions: governing equations
System (40)-(41) when all modes $\boldsymbol{p}$ are in Sector A describes rotation-free interactions between 3D modes, excited at scale $l_{f}$, and an imposed large-scale mean flow $\boldsymbol{U}(y)$. This is due to the frequencies of 3D modes in this sector being too slow, so that $\Delta_{\boldsymbol{p}}^{+-}\approx 1$ and rotation has no effect on the QL dynamics at leading order. For these modes, the same dynamical system would be obtained if $\Omega$ were put to 0 in Eq. (27).
@@
However, the helical basis does not diagonalize the stationary system of correlators $\Phi_{\boldsymbol{p}}^{s},\Psi_{\boldsymbol{p}}^{+-}$. Instead, the system can be triangularized, and possibly solved analytically, when formulating the fluid equations in terms of primary variables, as is classically done in mean-fluctuation descriptions of shear flows [42, 20]. In this section we will present the solution for the energy flux and Reynolds stress for this non-rotating, directly forced case. This solution will be useful for the treatment of modes in Sector A, which also exhibit such dynamics and can either be directly forced, or receive their energy through a flux in $p$ space, coming from modes in the homochiral sector.
@@
We consider a non-rotating system of mean flow $\boldsymbol{U}=U(y)\mathbf{e}_{x}$ with a given mean shear rate $U^{\prime}$}, and 3D perturbations $[u,v,w](x,y,z)$ such that $\boldsymbol{u}\cdot\nabla\boldsymbol{u}\ll\boldsymbol{u}\cdot\nabla\boldsymbol{U}$, which obeys the Rayleigh equations
$\partial_{t}\begin{bmatrix}\nabla^{2}v\\ \eta\end{bmatrix}=\begin{bmatrix}-(U\nabla^{2}-\partial_{y}^{2}U)\partial_{x}&0\\ -{U^{\prime}}\partial_{z}&-U\partial_{x}\end{bmatrix}\begin{bmatrix}v\\ \eta\end{bmatrix}+\begin{bmatrix}\nabla^{2}f_{y}\\ (\nabla\times\boldsymbol{f})_{y}\end{bmatrix}$ (63)
@@
with $\eta\equiv(\nabla\times\boldsymbol{u})_{y}$ the $y$-component of vorticity (see e.g [20]). Restricting Eq. (63) to our sinusoidal mean flow $U(y)={U^{\prime}}L_{y}/(\sqrt{2}\pi)\sin(2\pi y/L_{y}),U_{\boldsymbol{k}}=\frac{U^{\prime}}{i\sqrt{2}k_{y}}$ and using scale separation $L_{y}k_{f}\gg 1$ (consequently, $\partial_{y}^{2}U\ll U\nabla^{2}$), we obtain the simplified system in Fourier space:
$\partial_{t}(p^{2}v_{\boldsymbol{p}})$ $=\frac{p_{x}{U^{\prime}}}{\sqrt{2}k_{y}}\Bigg[|-\boldsymbol{p}-\boldsymbol{k}|^{2}v_{-\boldsymbol{p}-\boldsymbol{k}}^{*}-|-\boldsymbol{p}+\boldsymbol{k}|^{2}v_{-\boldsymbol{p}+\boldsymbol{k}}^{*}\Bigg]+p^{2}f_{\boldsymbol{p},y}$ (64)
$\partial_{t}\eta_{\boldsymbol{p}}$ $=\frac{{U^{\prime}}}{\sqrt{2}}\Bigg[\frac{p_{x}}{k_{y}}(\eta_{-\boldsymbol{p}-\boldsymbol{k}}^{*}-\eta_{-\boldsymbol{p}+\boldsymbol{k}}^{*})-ip_{z}(v_{-\boldsymbol{p}-\boldsymbol{k}}^{*}+v_{-\boldsymbol{p}+\boldsymbol{k}}^{*})\Bigg]+g_{y}$ (65)
with $k_{y}=2\pi/L_{y}$ and $g_{y}=ip_{z}f_{\boldsymbol{p},x}-ip_{x}f_{\boldsymbol{p},z}$ and where we have used that $p_{x}=-q_{x},p_{y}=-q_{y}\pm k_{y},p_{z}=-q_{z}$.
@@
We now write the equation for the ensemble-averaged cross-flow energy $\langle|v_{\boldsymbol{p}}|^{2}\rangle$, velocity-vorticity correlation $\langle\eta_{\boldsymbol{p}}v_{-\boldsymbol{p}}\rangle$ and cross-flow enstrophy $\langle|\eta_{\boldsymbol{p}}|^{2}\rangle$ at $O(k_{y}/p)$:
@@
$\partial_{t}\frac{\langle|v_{\boldsymbol{p}}|^{2}\rangle}{2}$ $=\frac{{U^{\prime}}}{\sqrt{2}}\Big[p_{x}\partial_{p_{y}}+\frac{4p_{x}p_{y}}{p^{2}}\Big]\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}+\epsilon\chi_{\boldsymbol{p},y}$ (66)
@@
$\partial_{t}\frac{\langle\eta_{\boldsymbol{p}}v_{-\boldsymbol{p}}\rangle}{2}$ $=\frac{{U^{\prime}}}{\sqrt{2}}\Bigg(\Big[p_{x}\partial_{p_{y}}+2\frac{p_{x}p_{y}}{p^{2}}\Big]\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}-ip_{z}\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}\Bigg)$ (67)
@@
$\partial_{t}\frac{\langle|\eta_{\boldsymbol{p}}|^{2}\rangle}{2}$ $=\frac{{U^{\prime}}}{\sqrt{2}}\Big(p_{x}\partial_{p_{y}}\langle\eta\eta\rangle_{\boldsymbol{p}\boldsymbol{k}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-}2p_{z}\Im\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}\Big)+\epsilon p^{2}\chi_{\boldsymbol{p},y}$ (68)
@@
$\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}\equiv\frac{1}{2}\Big(\langle v_{\boldsymbol{p}}v_{-\boldsymbol{p}+\boldsymbol{k}}\rangle+\langle v_{\boldsymbol{p}}v_{-\boldsymbol{p}-\boldsymbol{k}}\rangle\Big),~~~~~\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}\equiv\frac{1}{2}\Big(\langle\eta_{\boldsymbol{p}}v_{-\boldsymbol{p}+\boldsymbol{k}}\rangle+\langle\eta_{\boldsymbol{p}}v_{-\boldsymbol{p}-\boldsymbol{k}}\rangle\Big),~~~~~\langle\eta\eta\rangle_{\boldsymbol{p}\boldsymbol{k}}\equiv\frac{1}{2}\Big(\langle\eta_{\boldsymbol{p}}\eta_{-\boldsymbol{p}+\boldsymbol{k}}\rangle+\langle\eta_{\boldsymbol{p}}\eta_{-\boldsymbol{p}-\boldsymbol{k}}\rangle\Big)$ (69)
and used that $\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{*}=\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ since $\langle v_{\boldsymbol{p}}v_{-\boldsymbol{p}+\boldsymbol{k}}\rangle=\langle v_{-\boldsymbol{p}}v_{\boldsymbol{p}+\boldsymbol{k}}\rangle+O(k_{y}/p)$, as well as
$\langle v_{\boldsymbol{p}}v_{-\boldsymbol{p}-\boldsymbol{k}}\rangle$ $=\langle v_{-\boldsymbol{p}}v_{\boldsymbol{p}-\boldsymbol{k}}\rangle+O(k_{y}/p),~~~~~~~\langle v_{\boldsymbol{p}}v_{-\boldsymbol{p}-\boldsymbol{k}}\rangle=\langle v_{-\boldsymbol{p}}v_{\boldsymbol{p}-\boldsymbol{k}}\rangle+O(k_{y}/p),~~~~~~~\langle\eta_{\boldsymbol{p}}v_{-\boldsymbol{p}+\boldsymbol{k}}\rangle=\langle\eta_{\boldsymbol{p}-\boldsymbol{k}}v_{-\boldsymbol{p}}\rangle+O(k_{y}/p),$
$k_{y}\partial_{p_{y}}\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}$ $=\langle\eta_{p+k}v_{-p}\rangle-\langle\eta_{p}v_{-p+k}\rangle+O(k_{y}/p)=\langle\eta_{p}v_{-p-k}\rangle-\langle\eta_{p-k}v_{-p}\rangle+O(k_{y}/p),$ (70)